OpenStudy (anonymous):

1/4x - 1/5 (x-4)= 5x looking for where to start

hero (hero):

Hint: \[\frac{1}{4}x - 5x = \frac{1}{5}(x-4)\]

hero (hero):

\[\frac{1}{4}x - 5x = \frac{1}{5}x - \frac{4}{5}\]

hero (hero):

\[\frac{1}{4}x - 5x - \frac{1}{5}x = - \frac{4}{5}\] \[\left(\frac{1}{4} - 5 - \frac{1}{5}\right)x = - \frac{4}{5}\]

OpenStudy (anonymous):

ok

hero (hero):

\[-\frac{99}{20}x = - \frac{4}{5}\] \[ \frac{4}{5} =\frac{99}{20}x\]

hero (hero):

I hope that (x - 4) term wasn't in the denominator of the fraction.

OpenStudy (anonymous):

no, it wasnt

hero (hero):

Okay, so then, you should be able to figure things out from there.

OpenStudy (anonymous):

unfortunately, that is where I got stuck

OpenStudy (anonymous):

i ended up with -83/20 = x

OpenStudy (anonymous):

which was wrong

hero (hero):

Well of course. x isn't negative.

hero (hero):

I tried to show that in my very last entry.

OpenStudy (anonymous):

lol, it can be in my math world..... my daughter tells me the math police are going to arrest me for breaking math laws...

hero (hero):

But 88/20 isn't the right value anyway.

hero (hero):

And even if it was, you'd still have to reduce it to lowest terms.

OpenStudy (anonymous):

no, 83

OpenStudy (anonymous):

4/5 = 16/20

OpenStudy (anonymous):

and when you subtract the 99/20 from the right and do the same to the left, I came up with -83/20

hero (hero):

You don't subtract 99/20 from both sides.

hero (hero):

You have to multiply both sides by 20/99 to isolate x. Then afterwards, reduce.

OpenStudy (anonymous):

awwww and come up with 16/99!

hero (hero):

Congrats, you found the correct answer.

OpenStudy (anonymous):

with much in-trepidation